OpenStudy (anonymous):

The function f(x) is shown below: f(x) = (5)x - 1 The function f(x) is shifted to the right by 13 units. Which of the following best represents the new function? f(x) = (5)x + 13 f(x) = (5)x + 12 f(x) = (5)x - 14 f(x) = (5)x - 13

OpenStudy (anonymous):

I think it's f(x) = (5)^x - 13

OpenStudy (anonymous):

I'm not that sure though.

OpenStudy (anonymous):

Here's a screenshot of the question if that makes things easier.

OpenStudy (anonymous):

im pretty sure its b

OpenStudy (anonymous):

if you're shifting it to the right youre adding. -1+13=12

zepdrix (zepdrix):

Remember that when we make `horizontal` shifts, we do the opposite operation, yes? :) So to shift something to the right by 13, we would subtract 13 from our x.

zepdrix (zepdrix):

\[\Large\rm f(x)=5^{\color{orangered}{x}-1}\qquad\to\qquad F(x)=5^{\color{orangered}{x-13}-1}\]

OpenStudy (anonymous):

So it would be f(x) = (5)^x - 13? or f(x) = (5)^x - 14?

OpenStudy (anonymous):

since there's the \[5^x-13-1\]

zepdrix (zepdrix):

yes c: good.

OpenStudy (anonymous):

To be clear (sorry, I'm a complete dweeb when it comes to math), it'd be (5)^x - 14?

zepdrix (zepdrix):

Yes. 5^(x-14)

OpenStudy (anonymous):

Thank you so much, I appreciate it! :)

zepdrix (zepdrix):

np dweeb c; hah